flying against the wind a jet travels 4160 miles in 8 hours. flying with the wind, the same jet travels 3920 miles in 4 hours. what is the rate of the jet in still air and what is the rate of the wind
Answer by Boreal(15235)

mph against the wind is 4160/8=520 mph
mph with the wind is 3920/4=980 mph
x=plane speed
y=wind speed
x+y=980
x-y=520
2x=1500
x=750 mph, plane's speed
wind is 230 mph.